mathslover (mathslover):

Mathematics Quiz #2 : Get ready. Total Questions :10 Marks for correct answer : 4 Marks for wrong answer : -1 The first who answers will be awarded with +4. The second who answers with a short hint of the soln will also be awarded +4. Only 2 answers will be considered. Are you guys ready?

terenzreignz (terenzreignz):

Why the change of rules?

mathslover (mathslover):

It was required. :)

mathslover (mathslover):

2 minutes more. Then we will have the quiz started. Have any questions?

terenzreignz (terenzreignz):

Yes. What is the answer to the ultimate question of life, the universe, and everything?

OpenStudy (johnweldon1993):

42 :D

mathslover (mathslover):

Lol , only related to quiz :D

mathslover (mathslover):

@Zarkon @Mertsj @jim_thompson5910 @dumbcow @ganeshie8 [Test ur skills now :) ]

mathslover (mathslover):

Here comes first question. Good Luck all Let \(a_1 , a_2 , ... ,a _{10} \) be in AP and \(h_1 , h_2 , ... , h_{10} \) be in HP If \(a_1 = h_1 = 2\) and \(a_{10} = h_{10} = 3 \) then \(a_4 h_7 = ?\)

OpenStudy (anonymous):

6.48? lol

mathslover (mathslover):

@chris00 : -1

OpenStudy (anonymous):

o well.

mathslover (mathslover):

HP : Harmonic Progression GP : Geometric Progression AP : Arithmetic Progression

OpenStudy (johnweldon1993):

Wouldnt it be 6?

OpenStudy (anonymous):

totally did not know that...haha

mathslover (mathslover):

@johnweldon1993 : +4 One more chance left guys. But now provide soln

mathslover (mathslover):

(Provide hint) not soln...

OpenStudy (anonymous):

on top of it john! haah

mathslover (mathslover):

Come on friends ! Its not that difficult ;)

mathslover (mathslover):

Should we move on for next question ?

mathslover (mathslover):

2 minutes more !

mathslover (mathslover):

Time up !!!

mathslover (mathslover):

OK So here is the next question (Very very easy): If H is the harmonic mean between P and Q , then the value of (H / P) + (H / Q) is (in terms of P and Q)

mathslover (mathslover):

You will get 5 minutes for this question. Time starts now

OpenStudy (anonymous):

PQ/P+Q

mathslover (mathslover):

Surre

mathslover (mathslover):

@chris00 : -1

OpenStudy (anonymous):

i got no clue if the questions are similar. haha i'm out.

mathslover (mathslover):

@SharpEdge I need answer. Whats your answer ? @oldrin.bataku Same for u

terenzreignz (terenzreignz):

2

mathslover (mathslover):

@chris00 No problem. Keep trying

mathslover (mathslover):

@terenzreignz +4

mathslover (mathslover):

good work

mathslover (mathslover):

2 minutes more

mathslover (mathslover):

Solution is needed when you have posted the answer :)

OpenStudy (anonymous):

$$H=2/(1/P+1/Q)=2PQ/(P+Q)\\H/P=2Q/(P+Q)\\H/Q=2P/(P+Q)\\H/P+H/Q=2(P+Q)/(P+Q)=2$$

mathslover (mathslover):

@oldrin.bataku +4

mathslover (mathslover):

@SharpEdge +4 (your sol. was correct)

mathslover (mathslover):

Time up ! Next question: Another easy one : If the sum of the series 2, 5, 8, 11, …… is 60100, then n is

mathslover (mathslover):

You have got 8 minutes

mathslover (mathslover):

Good work @SharpEdge +4

OpenStudy (anonymous):

lol does that mean he gets +8?

terenzreignz (terenzreignz):

How gracing. It's (apparently) an arithmetic progression so \[\Large 60100 = \frac{n(a_1+a_n)}{2}=\frac{n(2+3(n-1))}{2}\]

OpenStudy (anonymous):

Clearly \(a_n=3n-1\) so:$$S_n=\frac{n(2+3n-1)}2=60100\\3n^2+n-120200=0\\n=200 $$

mathslover (mathslover):

@oldrin.bataku , @terenzreignz +4

terenzreignz (terenzreignz):

200, but it'd be redundant, as already mentioned :( oh well moving on :D

mathslover (mathslover):

Now wait for points table

terenzreignz (terenzreignz):

haha fail ^

terenzreignz (terenzreignz):

LOL jk

OpenStudy (anonymous):

Aren't you basically stealing attention from all the other posts? :-p now they'll never get answered!

terenzreignz (terenzreignz):

hush @oldrin.bataku you'll expose @mathslover 's grand plan!!!! haha jk

mathslover (mathslover):

@terenzreignz +8 @SharpEdge +8 @oldrin.bataku +8 @johnweldon1993 +4 @chris00 -2

OpenStudy (anonymous):

just u wait guys

OpenStudy (anonymous):

haha is that a sympathy medal?

terenzreignz (terenzreignz):

My Teamwork score went up a point.

mathslover (mathslover):

Next question : If the equation : \(x^4 - 4x^3 + ax^2 + bx + 1 =0\) has four positive roots, then (a,b) = ?

OpenStudy (anonymous):

Are they distinct roots?

mathslover (mathslover):

Not given.

OpenStudy (anonymous):

(-2,4)

OpenStudy (anonymous):

just a ping

mathslover (mathslover):

@chris00 , No :( -1

OpenStudy (anonymous):

(6,-4)

mathslover (mathslover):

@chris00 : +4 @SharpEdge : -1 :(

OpenStudy (anonymous):

BOOM

mathslover (mathslover):

One more chance left... Any one?

OpenStudy (anonymous):

$$x_1x_2x_3x_4=1\\x_1x_2x_3+x_2x_3x_4+x_1x_3x_4+x_1x_2x_4=b\\x_1x_2+x_2x_3+x_1x_3+x_2x_4+x_3x_4+x_1x_4=a\\x_1+x_2+x_3+x_4=4$$So clearly we have \(x_1=x_2=x_3=x_4=1\) yielding \(b=4\) and \(a=6\) yielding \((6,-4)\)

mathslover (mathslover):

Here it comes... @oldrin.bataku +4

OpenStudy (anonymous):

\((6,4)\)*

OpenStudy (anonymous):

sorry wrote it wrong I believe

mathslover (mathslover):

@oldrin.bataku ? 6,4 or 6,-4 ?

OpenStudy (anonymous):

ERG it should be \((6,-4)\)

OpenStudy (anonymous):

i kept flipping the signs when writing

mathslover (mathslover):

Ok. @oldrin.bataku : -1 and @SharpEdge : -1 @chris: +3

terenzreignz (terenzreignz):

@SharpEdge y so competitive? LOL

mathslover (mathslover):

Hmm... @oldrin.bataku : 2 and @SharpEdge : 2 because ur hints were right...

OpenStudy (anonymous):

hahahaha

terenzreignz (terenzreignz):

@mathslover your point system is so indecisive XD

mathslover (mathslover):

lol but thy were good

OpenStudy (anonymous):

$$x_1x_2x_3x_4=1\\x_1x_2x_3+x_2x_3x_4+x_1x_3x_4+x_1x_2x_4=-b\\x_1x_2+x_2x_3+x_1x_3+x_2x_4+x_3x_4+x_1x_4=a\\x_1+x_2+x_3+x_4=4$$We clearly see \(x_i=1\) and thus \(b=-4,a=6\) so \((6,-4)\)

mathslover (mathslover):

I can't neglect their work...

terenzreignz (terenzreignz):

If you say so...

mathslover (mathslover):

Support a bit guys... you all have some chances left. 6 quest. left

OpenStudy (anonymous):

COMING BACK FROM THE DEAD!

mathslover (mathslover):

Points table : @oldrin.bataku : 10 @SharpEdge : 10 @terenzreignz : 8 @chris00 : 1

mathslover (mathslover):

Oh and @johnweldon1993 : 4

mathslover (mathslover):

Find the maximum and minimum values of f(x) = x^2 - 5x + 6 [Question no. 5]

terenzreignz (terenzreignz):

No maximum -1/4 min

OpenStudy (anonymous):

No maximum but a minimum \((5/2,-1/4)\)

OpenStudy (anonymous):

we can do this by differentiating f(x) HINT

mathslover (mathslover):

@terenzreignz : + 4 @oldrin.bataku yes at x = 5/2 it is -1/4 +4 for u too

OpenStudy (anonymous):

$$f'(x)=2x-5\\2x=5\\x=5/2\\f(5/2)=25/4-25/2+6=-1/4$$

OpenStudy (anonymous):

and finding stationary points

mathslover (mathslover):

Points table : @oldrin.bataku : 14 @terenzreignz : 12 @SharpEdge : 10 @johnweldon1993 : +4 @chris00 : 1

mathslover (mathslover):

Be ready for next quest.